OpenStudy (anonymous):

Find a formula for the nth derivative of f(x) = e^(-x/3)

OpenStudy (anonymous):

you will need a -1^n

OpenStudy (anonymous):

(-1/3)^n . f(x)

OpenStudy (anonymous):

\[f ^{n}(x)=(\frac{-1}{3})^ne^{\frac{-x}{3}}\]

OpenStudy (anonymous):

Oh yeah, what he said :P

OpenStudy (anonymous):

(-1/3)^n . f(x)...does that always work?? replacing 1/3 with whatever is in the prob..

OpenStudy (anonymous):

no its only fr this case..

OpenStudy (anonymous):

In general:\[f ^{n}(e^m)=m^n e^m\]

OpenStudy (anonymous):

the general case is \[f^{(n)}(x) = (g'(x))^{n} . f(x) \] if \[f(x) = e^{g(x)}\]

OpenStudy (anonymous):

that helps alot! thanks guys

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

No problem.